The ascending arrangement of (2/3), (6/7), (13/21) is:
(a) 6/7, 2/3, 13/21 (b) 13/21, 2/3, 6/7 (c) 6/7, 13/21, 2/3 (d) 2/3, 6/7, 13/21

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (b) is correct here
LCM of 21, 3, 7 = 21
Now, let us change each of the given fraction into an equivalent fraction having 21 as the denominator.

[(13/21) × (1/1)] = (13/21) [(2/3) × (7/7)] = (14/21) [(6/7) × (3/3)] = (18/21)

Clearly,
(13/21) < (14/21) < (18/21)

Hence,
(13/21) < (2/3) < (6/7)

Hence, the given fractions in ascending order are (13/21), (2/3), (6/7)


Thanks